### v2

@@ Appendix B Details of Example 1
$z_{{\text{LP}}_{1}}=\min_{\bm{x}}~{}$ $x_{1}^{1}+x_{2}^{1}$ (57)
@@
$z_{{\text{LP}}_{2}}=\min_{\bm{x}}~{}$ $x_{1}^{1}+x_{2}^{1}$ (58)
@@
It is simple to check that $\bar{\bm{x}}$ is feasible to problem (57), and thus $z_{{\text{LP}}_{1}}\leq\frac{1}{6}$. It remains to show $x_{1}^{1}+x_{2}^{1}\geq\frac{1}{6}$ for any feasible solution $\bm{x}$ of (57). By carefully combining (50), (52), and (53), we have
@@
It is simple to check that $\hat{\bm{x}}$ is feasible to problem (58), and thus $z_{{\text{LP}}_{2}}\leq\frac{1}{4}$. It remains to show $x_{1}^{1}+x_{2}^{1}\geq\frac{1}{4}$ for any feasible solution $\bm{x}$ of (58). By (50) and (56), we have $3\geq 2x_{3}^{1}+2x_{3}^{2}\geq 2x_{3}^{1}+2x_{3}^{1}=4x_{3}^{1}$ and thus $x_{3}^{1}\leq\frac{3}{4}$, which, together with (48), implies $x_{1}^{1}+x_{2}^{1}\geq\frac{1}{4}$.

### v3

@@ Appendix B Details of Example 1
$z_{{\text{LP}}_{1}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname*{minimize}_{\bm{x}}}}~{}$ $x_{1}^{1}+x_{2}^{1}$
@@
$z_{{\text{LP}}_{2}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\operatorname*{minimize}_{\bm{x}}}}~{}$ $x_{1}^{1}+x_{2}^{1}$
@@
It is simple to check that $\bar{\bm{x}}$ is feasible to problem (62), and thus $z_{{\text{LP}}_{1}}\leq\frac{1}{6}$. It remains to show $x_{1}^{1}+x_{2}^{1}\geq\frac{1}{6}$ for any feasible solution $\bm{x}$ of problem (62). By carefully combining (55), (57), and (58), we have
@@
It is simple to check that $\hat{\bm{x}}$ is feasible to problem (63), and thus $z_{{\text{LP}}_{2}}\leq\frac{1}{4}$. It remains to show $x_{1}^{1}+x_{2}^{1}\geq\frac{1}{4}$ for any feasible solution $\bm{x}$ of problem (63). By (55) and (61), we have $3\geq 2x_{3}^{1}+2x_{3}^{2}\geq 2x_{3}^{1}+2x_{3}^{1}=4x_{3}^{1}$ and thus $x_{3}^{1}\leq\frac{3}{4}$, which, together with (53), implies $x_{1}^{1}+x_{2}^{1}\geq\frac{1}{4}$.